### v1

Abstract
While sentiment analysis has advanced from sentence to aspect-level, i.e., the identification of concrete terms related to a sentiment, the equivalent field of Aspect-based Emotion Analysis (ABEA)is faced with dataset bottlenecks and the increased complexity of emotion classes in contrast to binary sentiments. This paper addresses these gaps, by generating a first ABEA training dataset, consisting of 2,621 English Tweets, and fine-tuning a BERT-based model for the ABEA sub-tasks of Aspect Term Extraction (ATE)and Aspect Emotion Classification (AEC).
The dataset annotation process was based on the hierarchical emotion theory by Shaver et al. [1] and made use of group annotation and majority voting strategies to facilitate label consistency. The resulting dataset contained aspect-level emotion labels for Anger, Sadness, Happiness, Fear, and a None class. Using the new ABEA training dataset, the state-of-the-art ABSA model GRACE by Luo et al. [2] was fine-tuned for ABEA. The results reflected a performance plateau at an F1-score of 70.1% for ATE and 46.9% for joint ATE and AEC extraction. The limiting factors for model performance were broadly identified as the small training dataset size coupled with the increased task complexity, causing model overfitting and limited abilities to generalize well on new data.
keywords:
aspect-based, emotion analysis, GRACE, hyperparameter optimization, data annotation
1 Introduction
One of the largest and most widely researched sub-fields of Natural Language Processing (NLP) is Sentiment Analysis (SA). In its most basic form, SA assigns sentiment polarity labels (positive, neutral, and negative) to systematically describe the opinions expressed in text [3, 4]. Beyond basic polarity classification, SA can encompass several other classification tasks, including sarcasm or subjectivity detection [4]. SA can be categorized into different levels based on the scope of the analysis. Document-level SA assigns a single sentiment to an entire document (e.g., article), while sentence-level SA evaluates individual sentences [5]. Recent advancements in computing resources, model architectures, and training data availability have allowed SA research to progress toward a more fine-grained sub-sentence level, so-called aspect-level analysis [6, 7, 8]. It can generate individual outputs for every word or token, where tokens may even be sub-word units [9]. Aspect-based Sentiment Analysis (ABSA) can involve one or more subtasks, which are either addressed sequentially, i.e., individually and then combined to draw conclusions, or jointly as simultaneous outputs of a model [10].
Emotion Analysis (EA) is a sub-discipline of SA that categorizes texts according to distinct emotion classes instead of sentiment polarities. In contrast to SA, the field of EA has yet to undergo a comparable transition to such detailed, sub-sentence level analyses [11, 12, 13, 14]. Given the lack of Aspect-based Emotion Analysis (ABEA) concepts and taxonomies, we follow Zhang et al. [15] for the definitions of the following two subtasks of ABEA (cf. Figure 1): 1) Aspect Term Extraction (ATE): finding the target word(s) of the sentiment in the text, 2) Aspect Emotion Classification (AEC): assigning an emotion label to the respective aspect. In these definitions, an aspect is understood as the target word(s) of an emotion in a given text.
Figure 1: Two sub-tasks of ABEA (adapted from Zhang et al. [15]).
The value of large-scale social media analyses is widely recognized for its direct link to citizen opinions across spatio-temporal dimensions. In ABSA, several studies have incorporated social media data, in particular Twitter data. However, no equivalent studies have emerged thus far for ABEA, leaving a wide research gap with much potential for new insights, methods, and datasets. This gap is particularly relevant for domains such as disaster management, where social media data is increasingly used for real-time monitoring, situational awareness, and response coordination by gauging public needs and opinions [16, 17, 18].
Even studies reviewing the more prominent field of ABSA have highlighted significant dataset-related challenges, such as domain specificity restraining the generalizability of models [19, 20, 10, 21, 22]. Filling the training data gap for ABEA, particularly for social media data, is a prerequisite for any successful model fine-tuning and application. This paper discusses the logistical challenges of generating such a new ABEA training dataset, which encompass the appropriate selection and integration of emotion theory, accounting for human subjectivity [23], and assessing inter-annotator agreement (IAA) at the sub-sentence level. Our second research aim is to implement a model capable of ABEA through adapting the Gradient-Harmonized and Cascaded Labeling for Aspect-Based Sentiment Analysis (GRACE) model by Luo et al. [2]. This endeavor presents various sub-challenges, including the choice of model hyperparameters and methods for performance validation. Consequently, we aim to address the following research questions:
  • RQ1. How can a consistent training dataset for Aspect-based Emotion Analysis be annotated from social media data?
  • RQ2. What is the efficacy of repurposing an Aspect-based Sentiment Analysis model for Aspect-based Emotion Analysis? Which hyperparameter configurations significantly affect model performance?
2 Related works
2.1 Theories of Emotion
An important basis for EA is conceptualizing how emotions are expressed, categorized, and manifested [24, 11, 25, 26]. However, there is no clear consensus on this subject matter, making direct comparisons between research outputs challenging [24, 27]. Originating from the discipline of psychology, two main types of emotion models are commonly used in EA: Categorical models, such as Ekman [28], classify emotions into distinct groups, whereas dimensional models, such as Plutchik [29], position them on continuous scales, such as happiness-sadness or calmness-excitement. Shaver et al. [1] conceptualize a hierarchical emotion graph that incorporates elements of both categorical and dimensional approaches. They identify a set of basic emotions (Love, Joy, Anger, Sadness, Fear, optionally Surprise), which serve as prototypes from which more complex emotions branch out. The closer any two emotion clusters are positioned along the x-axis, the more they are related. Higher y-axis splits indicate stronger sub-branch relationships. This model acknowledges that emotions are not entirely distinct categories but exist on a continuum, with variations within each basic emotion. The hierarchical emotion model [1] therefore forms the theoretical basis for the ABEA model in this study.
2.2 Training datasets for ABSA & ABEA
Alongside fine-tuning methods, an equally defining factor for the adaptation of pre-trained models for downstream tasks is the availability of task-specific training data. Current research on ABSA is limited by the number of publicly available training datasets and their domain specificity. They are mostly based on customer reviews, which are sourced from e-commerce platforms such as Amazon or Yelp [30]. The most widely used benchmark datasets “SemEval” [31, 32, 33] consist of approximately 3,000 restaurant and laptop reviews and are labeled with aspect terms and sentiment polarity. Alternative ABSA training datasets include “SentiHood” [22], which contains 5,215 sentences with one or more sentiment target terms related to urban neighborhoods, and “MAMS” [34], where each sentence includes multiple aspects and sentiments.
For social media analyses, well-structured datasets like SemEval are unsuitable since they do not resemble the typical complexity, noise, and unstructured nature of social media texts. Three ABSA training datasets consisting of Twitter posts can be found in published literature [35, 36, 37], which are intentionally biased towards named entities like celebrities, companies, or well-known products. In addition, all predominantly contain neutral sentiment labels. Though further social media ABSA training datasets exist, they either do not extract aspect terms [38], or are structured in complex formats like dialogue [39].
Though various datasets for sentence-level emotions have been created [24, 13], ABEA research suffers from a lack of publicly available training datasets annotated with both aspect terms and their corresponding emotions [11] (cf. Table 7). Only the datasets in Maitama et al. [40], De Geyndt et al. [41], Üveges et al. [42] adhere to the definition of aspect terms as the target of an emotion. Among them, only one dataset is entirely in English, and none are publicly available. Furthermore, none of the datasets share the same emotion classification schema. It should also be noted that all datasets are domain-specific, being related to finance and stock markets [43] or Hungarian speeches [42], making them less applicable and effective for general-purpose fine-tuning of ABEA models [10].
2.3 Models for ABSA
So far, individual ABSA tasks have been more widely researched and have reported higher performance metrics compared to compound tasks [15]. This performance drop is likely due to the differing nature of the sub-tasks, which are difficult to integrate into a single model [2]. For example, ATE and Aspect Sentiment Classification (ASC) are inherently different NLP tasks, the first being a sequence labeling task and the latter being a classification task. Despite this added complexity, unified models offer several advantages over pipeline methodologies, such as their ability to learn co-dependencies and interactions between the individual tasks [2], while also being able to avoid error propagation [44].
The performance of both single-task and compound-task models is most often assessed against the SemEval datasets [31, 32, 33], with most publications reporting the macro F1-score as a means for comparison. In this, single-task ABSA models generally achieve higher performances than compound-task ABSA models. ATE-only and ASC-only models have reported F1-scores up to 89% and 91%, respectively [45, 46]. In comparison, compound ABSA methodologies lag behind in comparative performance, with pair-extraction models reaching F1-scores in the mid-70% to 80% range [47, 48, 49], the highest triplet-extraction models achieving an F1-score of 72% [50, 51, 52], and quad-extraction of 63% [53, 54, 55].
2.4 Models for ABEA
Many ABEA publications do not focus on or even perform ATE from the input text. Padme and Kulkarni [25] perform sentence-level EA, first identifying general topics using the topic modeling technique Latent Dirichlet Allocation (LDA) and then creating subsets for each topic through keyword filtering. A similar approach is seen in Suciati and Budi [26], where aspect terms are predefined categories and their extraction is not part of the methodology. Instead, they compare how well their methods classify emotions for sentences belonging to the predefined aspect categories. Likewise, Sirisha and Chandana [56] report the use of a hybrid ABSA-RoBERTa-LSTM model to classify Tweets into sentiment and emotion classes without specifying how aspect terms were identified. While Dehbozorgi and Mohandoss [19] seem to extract aspect terms from input text, their method begins by conducting what appears to be sentence-level emotion classification using the Text2Emotion Python package.
These examples highlight the fundamentally different task conceptualizations in the small field of ABEA research, making systematic method comparisons difficult. A common characteristic emerging from the literature is the widespread use of pipeline methods to address the individual subtasks separately. This shows that ABEA methodologies have not yet advanced to joint or unified methods, which are the current research frontier in ABSA.
3 Methodology
The methodology presented in this paper consists of two main steps, namely the creation of an ABEA training dataset and the fine-tuning of the GRACE model [2] for ABEA. Figure 2 provides a schematic, high-level overview of our methodology. The code used in this paper is available on GitHub.
Figure 2: Overview of methodology.
3.1 Data
Social media data from the microblogging platform Twitter (now X) was used to create a ABEA training dataset. A Java-based crawler was employed to collect data, including the Tweet text, a timestamp, and the geolocation, via the v1.1 recent search and streaming Application Programming Interfaces (APIs) [57, 58]. A Tweet’s geolocation was stored either in the form of a point or, more commonly, as a rectangular polygon encompassing a region tagged by the user [59].
Based on the spatial and temporal attributes of the data, an English training dataset consisting of 75% random and 25% disaster-related social media posts was assembled. For this, Tweets from the United Kingdom (UK) and the United States (US) were combined to capture variations in the English language. Tweets for the annotation were then randomly sampled from two larger datasets detailed in Table 1. We incorporated domain-specific training data, assuming that it would contain a higher amount of emotional Tweets and thus addressing the overall class imbalance. The Californian dataset was binary classified with “related” or “unrelated” labels using an XLM-RoBERTa-base model [60] which was fine-tuned specifically for the classification of disaster-relatedness in social media data [61]. Thus, the training dataset was slightly skewed towards disasters, specifically wildfires and floods, potentially increasing the model sensitivity towards emotion and target detection in natural disaster contexts.
Table 1: Overview of data used to create an ABEA training dataset.
Attribute Random Sample Disaster-Related Sample
Date Range 2020-09-01 to 2021-02-28 2022-01-01 to 2022-09-30
Geographic Range London, UK California, USA
Total Dataset Size 25,000 16,336
Ratio 75% 25%
Bounding Box (EPSG: 4326)
-126.2721, 31.5850;
-112.5449, 42.7965
-0.675965, 51.1500;
0.47999, 51.8297
3.2 Training dataset annotation
Due to the lack of publicly available social media ABEA training datasets, a new dataset was annotated (cf. Table 1). In total, seven human annotators from the fields of data science and geoinformatics were enlisted for this task. Each Tweet was labeled by three annotators using an online tool configured with Doccano [62]. The final number of Tweets annotated by three annotators was 2,766. In these Tweets, aspect terms were identified and marked with one of the four basic emotion categories: Happiness, Anger, Sadness, or Fear. For dataset consistency, a fifth, null category None was also included. The annotation followed an iteratively designed guideline, which set out the tools, rules, and processes for annotations. The guideline was tested on a small pilot dataset of 50 Tweets and revised prior to being used for the annotation.
3.2.1 Emotion detection
For the annotations, two amendments were made to the hierarchical emotion graph by Shaver et al. [1]: Love and Joy were joined into one emotion Happiness, following the suggestion that even Ekman’s basic emotions can be consolidated due to high similarity between certain pairs (e.g., Anger and Disgust) [63]. Furthermore, Surprise was excluded as a stand-alone emotion class as it is a more transient, secondary emotion that triggers another, more dominant one [64]. The adapted graph was provided to the human annotators to help their decision making process.
The inclusion of a neutral class None is common practice in EA to ensure consistency in the annotations, properly handle texts without apparent emotions [65, 66], and ultimately enhance model robustness. This approach reduces computational complexity in identifying aspect terms and prevents overfitting by providing a label for unemotional texts, mitigating bias toward inapplicable emotion classes.
Annotators were first guided on how to detect explicit and implicit emotional cues within the text. Explicit emotions are signaled clearly through emotional words, such as "love" in the phrase “I love this time of year”, or "scared" in the phrase “my dogs are scared of the fireworks”. Explicit emotional references typically allow for straightforward identification of the relevant basic emotion. By contrast, implicit emotions require interpretation of the context (e.g., “our government isn’t fit to run a race”) or contextual cues like emojis, punctuation (e.g., “…” or “!!!”), or specific word spellings (e.g., “look at thaaaaaat”) to infer the respective emotion. Additionally, annotators were cautioned that some Tweets may seem emotionally charged but are merely factual reports of events without any intended emotional tone.
To keep the emotion annotations as objective as possible, the guideline addressed the risk of personal bias. Annotators were warned that their personal opinions, mood, or recent annotation experiences could affect their interpretation of a Tweet’s emotional content. For example, if an annotator recently encountered an angry Tweet about wildfires, they might be predisposed to interpret a subsequent Tweet on a related topic as expressing anger, even if the emotion is more subtle or different. To mitigate this, annotators were encouraged to consider the Tweet in various contexts and test different readings before settling on an emotion.
Lastly, sarcasm was highlighted as a particularly complicated part of emotion detection as it conveys an emotion opposite to the literal meaning of the words. For instance, a sarcastic Tweet like "I lost my key, how wonderful" may literally convey joy but actually express anger or sadness. The annotators were advised to carefully consider the broader context and the potential for sarcasm when determining the true emotional content to ensure the labeled emotion accurately represents the intended sentiment rather than the superficial meaning of the words.
3.2.2 Aspect term detection
Once an annotator identified an emotion in a Tweet, the associated aspect term, i.e., the specific target within the Tweet that the emotion is directed towards, was labeled. In the case of the None class, the main subject matter of the Tweet was chosen as the aspect term. If no emotion target or main subject was present, no aspect term was annotated. An example of such a Tweet was “seems fun”.
Table 2: Examples of aspect terms. The respective aspect term is marked in bold.
Category Type Examples
Nouns Singular noun “The smoke is getting closer”
“my car just died on me.”
Plural noun “Just thinking of all the holidays I’ll go on next year.”
Proper noun “I miss Obama
Hyde Park is so good for relaxing after work.”
Pronouns Subjective personal pronoun “She lost her home.”
Objective personal pronoun “I feel sorry for them
“I love you.”
Demonstrative pronoun This is amazing!”
These are delicious.”
Indefinite pronoun “Stay safe everyone
“Well done all!”
Verbs “I hate voting
“Would rather be sleeping right now”
The decision to predominantly focus on nouns, pronouns, and verbs stemmed from the otherwise significant increase in complexity in the structure of the labeled dataset. The test pilot revealed that many emotion targets are in fact a highly specific and lengthy combination of entities or actions. For example, in the phrase “the democrats giving these speeches need to get a grip on the facts”, neither “democrats” nor “speeches” fully encapsulate the target of the author’s frustration, requiring the annotation of “democrats giving these speeches” as the aspect term. While this more loosely defined annotation may better reflect the “ground truth” of humanly identifiable emotion targets, it greatly diminishes any systematization of aspect terms, making it exceedingly difficult to set a universal definition. Preliminary fine-tuning tests showed that such a loosely defined aspect term definition hinders a model’s ability to learn generalizable ATE patterns. Consequently, the definition of aspect terms was tightened, encouraging the annotators to focus on nouns (singular, plural, or proper), pronouns (personal, demonstrative, or indefinite), or verbs (cf. Table 2). The annotators were instructed to be concise and mark only the essential word(s) directly liked to the emotion.
3.2.3 Inter-annotator agreement
Given the complexity of human language, a high degree of ambiguity related to the annotation of aspect terms was expected. To address the risk of low IAA upfront, the annotators were invited to participate in in-person annotation sessions, where they annotated independently but advanced through the dataset at the same pace, allowing for discussions and clarifications at every Tweet. This method was inspired by Luo et al. [10] and Maitama et al. [40], who employed similar collaborative annotation methods.
All final annotations were determined by means of majority vote [20, 26]. At the character level, the annotated aspect terms of all three annotators were compared and only sequences marked by at least two annotators were kept. If all three annotators agreed that no aspect terms were present, the Tweet was included in the final dataset without aspect term labels. In cases where all three annotators marked aspect terms but none of the labeled sequences overlapped, the Tweet was excluded from the final dataset. A similar method was applied for the final emotion labels, with the additional step that cases without a majority emotion (i.e., all different emotion labels) were manually reviewed. Using this method, 5,704 characters (11%) from the 51,962 initially annotated aspect term characters were removed from the dataset.
3.2.4 Dataset statistics
Following the IAA assessment, the final ABEA training dataset consisted of 2,621 Tweets. The dataset was split into training, validation, and testing subsets using a 70%, 10%, and 20% proportion, respectively. The most common emotion was Happiness, which was assigned to 2,153 (39%) of the aspect terms. By contrast, only 282 (5%) of all aspect terms were labeled with Fear. Figure 3 gives an overview of the label frequencies.
Figure 3: Emotion labels in the ABEA training dataset.
A breakdown of aspect terms by part of speech tags revealed that most were nouns. 45% of all aspect terms were singular nouns such as “fire”, 16% were singular proper nouns such as “London”, and 13% were plural nouns such as “days”. Aside from nouns, 6% of aspect term words were pronouns such as “you”, and 5% were adjectives such as “relevant” (cf. Figure 7). Compared to the ABSA datasets by Mitchell et al. [35] and Dong et al. [36], this new dataset was both less domain-bound and had more diverse part of speech tags. While all emotion classes shared some aspect terms (e.g., “you”, “it”, “fire”), there were words unique to each emotion class, such as "shit" for Anger, "behavior" for Fear, "end" for Sadness, "today" for Happiness, and "month" for None. In the ABEA dataset, aspect terms ranged from 2 to 58 characters, with an average of 10. In comparison, aspect terms in the dataset by Mitchell et al. [35] ranged from 3 to 84 characters, averaging 11.
3.3 ABEA model fine-tuning
3.3.1 GRACE: Gradient-Harmonized and Cascaded Labeling for Aspect-Based Sentiment Analysis
With an annotated training dataset in place, a suitable model architecture was needed as a starting point for fine-tuning. In the absence of ABEA models, ABSA models were considered. Specifically, the search was narrowed down to models that jointly perform the two End-2-End ABSA tasks (ATE and ASC) and were reportedly adapted to social media data.
The applicable models are shown in Table 3 and were applied to the Twitter dataset by Mitchell et al. [35]. Most models reported considerably higher performance for the individual sub-tasks compared to compound ATE and ASC. For instance, the GRACE model [2] achieved 75.7% on ATE, while SpanABSA [67] achieved 75.3% and 75.2% for ATE and ASC, respectively. GRACE performed best on Twitter data, with a reported F1-score of 58.3% for joint ATE and ASC and was therefore selected as starting point for ABEA fine-tuning.
Table 3: Performance Comparison of End-2-End ABSA models on social media data.
Paper Model Method F1-Score [%]
[67] SpanABSA BERT-based span-extraction and polarity classification 57.7
[68] Unified-LSTM Unified stacked LSTM with boundary detection and gating 48.0
[2] GRACE Gradient harmonized, cascaded labeling BERT-based model 58.3
[44] MTMVN Multitask Multiview network with CNNs and BiGRUs 47.9
[69] CasNSA Cascade BERT-based model with hierarchical structure 50.1
[70] BDEP-IAM Block-level dependency syntax parsing with interactive attention 58.0
The basic structure of GRACE, which is adapted from the 12-layer, BERT-base model [71], consists of two computational branches for ATE and ASC respectively, which share several root layers. Luo et al. [2] split Bidirectional Encoder Representations from Transformers (BERT) into two branches, sharing the bottom $l$ (9) layers while reserving the top  $l:L$ (3) layers for task-specific ATE and ASC learning. This architecture enables a shared foundational understanding across both branches. Instead of treating ASC as a classification task, both tasks are implemented as sequence labeling to help facilitate co-extraction in a single, unified model. The ATE branch labels tokens as $B$ (beginning-), $I$ (inside-), or $O$ (outside of an aspect term). These labels in turn serve as additional input for the ASC branch, where a modified Transformer decoder generates sentiment labels $POS$ (positive), $NEU$ (neutral), $NEG$ (negative), or $O$ (other words). The passing of ATE label outputs as ASC inputs forms a cascaded labeling approach, which Luo et al. [2] argue enhances the model’s sentiment judgment.
In addition to cascaded labeling, Luo et al. [2] enhance GRACE’s training with the use of Gradient Harmonized Loss Function (GHL) and Virtual Adversarial Training (VAT), both of which are designed to improve model robustness when given difficult or imbalanced training data. GHL, introduced by Li et al. [72], mitigates class imbalance by weighting training examples based on their difficulty according to the model’s predictions during training. This increases model focus on underrepresented, difficult aspect terms, while also preventing overprediction of easier and more frequent neutral tokens. VAT, further enhances model robustness by introducing small adversarial perturbations to the input embeddings during the training process, improving model resilience to data variations [2].
Using these customizations, the fine-tuning was conducted across three training steps, each with a specific number of epochs, learning rate, and setting for VAT and GHL. The first two training steps focused exclusively on ATE. The key difference was that VAT was applied in the second step, but not the first. The third step then incorporated the training of the ASC branch for the full co-extraction of ATE and ASC. Luo et al. [2] configured the hyperparameters based on the SemEval laptop and restaurant datasets. With the extensive customizations for both the model architecture and training process by Luo et al. [2], the GRACE model appeared well suited for the complexities of joint ATE and AEC tasks, learning both tasks in tandem and optimizing the training for robustness.
3.3.2 Hyperparameter optimization
The GRACE model was initialized in the same manner as described by Luo et al. [2]. Due to the increased classification complexity of emotion classes and considerable structural differences between the SemEval and the new social media ABEA dataset, extensive Hyperparameter Optimization (HPO) was required to identify optimal hyperparameter configurations for ABEA.
The ABEA dataset was split into train, validation, and test subsets, using the same 70%-10%-20% proportions as Luo et al. [2]. In addition to standard hyperparameter categories such as batch size, learning rate, and warm-up proportion, several architectural modifications were tested, including adapting the number of shared layers $l$ for the ATE and ASC branches, the number of decoder blocks in the ASC branch, the classification head architecture, and shallow layer freezing during fine-tuning. Across the three overall training steps, the hyperparameters for the learning rate, GHL-use, VAT-use and batch size were tested individually for each training step. Using the reported optimal ABSA hyperparameter configurations as a baseline [2], hyperparameter configurations were adjusted in both directions where possible. Table 5 gives an overview over selected model specifications.
The adjustment of architectural elements and training settings enable more influence over the model’s fundamental training and structure. First, GHL and VAT configurations were adjusted to evaluate their respective impact on performance. Second, the number of shared layers between the ATE and ASC branches were adjusted in both directions to test the impact of more and less shared learning. When fewer layers are shared, each task-related branch of the GRACE model learns more task-specific representations. Third, the amount of ASC fine-tuning was extended beyond the originally designated epochs of the third training step. Towards this end, fine-tuning iterations were conducted in which ASC was trained from the outset, from the second training step, and as an added fourth training step. In the last approach, all other model layers related to ATE were frozen to conduct independent emotion learning. Fourth, shallow layer freezing was tested by keeping the bottom three, five, and nine BERT layers of the GRACE model frozen for the entirety of the fine-tuning process to preserve the originally learned parameters from pre-training. Lastly, the model’s single linear classification layer was replaced with more complex classifiers. These included a three-layered multilayer perceptron with rectified linear unit activation functions and a 0.1 dropout rate, as well as three, four, and five layered dense classifiers inspired by the DenseNet structure [73], which use linear layers and concatenation of intermediate results to preserve input information while improving feature representation.
HPO was conducted in a manual grid-search manner, mainly due to the need for code adjustments between iterations and separate validation steps after each fine-tuning iteration. Throughout the HPO process, recall, precision and F1 scores were used as performance indicators for both ATE and AEC task. Accuracy was generally not considered as it includes true negatives (TN) in its calculation. At the aspect-level, there is a vast number of TN values, i.e., all words in a sentence that are not aspect terms. Their high proportion skews the accuracy metric, which was consistently above 90%, and does not give a meaningful reflection of model performance.
Using these metrics, the model’s performance was systematically assessed for each HPO iteration. A baseline comparison was created with the originally reported hyperparameters by Luo et al. [2]. For each hyperparameter category $H_{c}$ where $c\in~{}${$batch~{}sizes,learning~{}rates,...$}, fine-tuning was performed for each configuration $H_{C_{i}}$ where $i\in$$config_{1},config_{2},...config_{n}$}. The averaged performance $P(H_{C_{i}})$ for each configuration was computed from seven sets of macro F1-scores (Equation 1). Specifically, the F1-score was calculated for both ATE and joint ATE +AEC using the validation data $\mathbb{D}_{\text{val}}$ and after each of the three training steps using the test data $\mathbb{D}_{\text{test}}$ (Equation 2). After fine-tuning, additional F1-scores were calculated for ATE utilizing the external ABSA datasets SemEval Laptop $\mathbb{D}_{\text{L}}$ and SemEval Restaurant 2014-16 $\mathbb{D}_{\text{R}}$ [31, 32, 33]. For AEC, a similar validation was performed using the SemEval Affect 2018 dataset $\mathbb{D}_{\text{A}}$ [74]. The performance of each configuration was then assessed against the previous best result. If any configuration from the current hyperparameter category produced a higher average performance than the previously recorded best, that configuration was retained for all subsequent iterations of HPO. Otherwise, the configuration was discarded to maintain the original settings across subsequent iterations.
$P(H_{C_{i}})=\frac{1}{7}\big{(}P_{\text{train}}+P_{\text{extern}})$ (1)
$P_{\text{train}}=\sum_{d\in(\mathbb{D}_{\text{val}},\mathbb{D}_{\text{test}})}\big{(}f_{1}^{d}(ATE,H_{C_{i}})+f_{1}^{d}(ATE+AEC,H_{C_{i}}))$ (2)
$P_{\text{extern}}=f_{1}^{\mathbb{D}_{R}}(ATE,H_{C_{i}})+f_{1}^{\mathbb{D}_{L}}(ATE,H_{C_{i}})+f_{1}^{\mathbb{D}_{A}}(AEC,H_{C_{i}})$ (3)
For validating the AEC task, the SemEval Affect 2018 dataset [74] was used, which consists of sentence-level emotion labels (joy, anger, sadness, fear), as no alternative ABEA validation datasets exist. When comparing the model’s aspect-level emotion inferences with these sentence-level labels, aspect-level emotions were aggregated at the sentence-level, excluding cases with multiple detected emotions per sentence to ensure consistency. To evaluate the co-extraction of ATE and AEC, precision, recall, and the F1-score were determined by focusing only on correctly identified aspect terms. A prediction was considered correct only if both the aspect and its emotion were accurately predicted.
4 Results
The configurations and results for selected hyperparameters, training, and model adjustments are shown in Table 5. It provides an overview of the model performance using F1-scores, for the new ABEA training dataset and for the validation tests performed on the $\mathbb{D}_{L}$, $\mathbb{D}_{R}$, and $\mathbb{D}_{A}$ datasets. The baseline (configuration 1) is given by the hyperparameters reported by Luo et al. [2] as the optimal settings for ABSA.
The highest performance - averaged across the ABEA validation, ABEA test, and SemEval external validation datasets - was achieved with the hyperparameter and model settings in configuration 41, with an overall average of 50.8%. Configuration 41 achieved slightly higher precision than recall for both the ATE-only task and the joint ATE and AEC tasks (see Table 4).
Table 4: Performance metrics on test data for configuration 41.
Configuration Recall [%] Precision [%] F1-Score [%]
ATE Only 68.8 71.5 70.1
ATE + AEC 46.1 47.7 46.9
Table 5: Hyperparameter Settings and Results. The table shows a selection of hyperparameter configurations with settings that were propagated onwards. Empty cells imply original settings as shown in Config. 1
Config. Nr. 1 6 12 14 33 41
Categories Orig Settings Batch Size Epochs Warmup AEC Layers VAT
Hyperparameter Settings
Dropout 0.1
Weight Decay 0.01
Batch Size 32 8
Gradient Acc. 2
Epochs Step 1 5 10
Epochs Step 2 3 9
Epochs Step 3 10 25
Warmup Method Linear Constant
Warmup Proportion 0.1
LR Training Step 1 3.0E-05
LR Training Step 2 1.0E-05
LR Training Step 3 3.0E-06
Shared Layers 9 5
Nr. of AEC Layers 2 6
VAT enabled in 2nd step disabled
F1 for the New ABEA Dataset [%]
ATE Validation 65.3 67.5 67.8 68.5 68.6 70.1
ATE + AEC Validation 42.0 41.4 42.5 43.6 44.7 46.9
ATE Test 65.9 67.0 68.2 67.2 66.8 67.1
ATE + AEC Test 42.9 45.1 44.7 43.6 45.9 47.1
ATE on $D_{R}$ 63.4 63.2 64.0 64.5 63.5 64.3
ATE on $D_{L}$ 46.5 46.1 46.8 47.4 45.8 46.2
AEC on $D_{A}$ 10.5 12.5 13.1 13.2 13.9 14.1
Average 48.1 49.0 49.6 49.8 46.5 50.8
Configurations 2–5, which adjust dropout and weight decay, as well as configurations 47-49, which implement layer freezing during training, were designed to address the likely tendency for overfitting of a complex BERT-based model when fine-tuned on a comparatively small dataset. During configurations 2–5, the initial low number of training epochs was maintained, while dropout was increased across all BERT layers in both task branches and the weight decay parameter in the loss function was increased to penalize large weights. The results, however, showed that in particular the performance on the SemEval validation datasets dropped to lower levels than those of the initial configurations. For configurations 47–49, which applied shallow layer freezing, no improvements were recorded for either the ABEA dataset metrics or the SemEval validation metrics.
Configurations 50–54, in which the classification heads were exchanged with slightly more complex architectures, exhibited the most prominent trade-offs between gains in one metric and declines in others. All configurations, including the baseline, consistently demonstrated low performance in the emotion validation on $\mathbb{D}_{A}$.
Hyperparameter adjustments that were ultimately kept included a reduction of the batch size (configuration 6), an increase in the number of training epochs (configuration 12), an adjustment to the warm-up method (configuration 14), a reduction in the number of shared layers between the ATE and AEC branches and an increase of AEC branch layers (configuration 33), and the omission of VAT in the second training step.
The fine-tuning results for the baseline (configuration 1) and the highest average configuration (configuration 41) were further validated through 10-fold cross-validation. The results are displayed in Table 6 and show a 6-point improvement for the joint ATE and AEC tasks over the baseline. Figure 4 shows the progression of model performance during the fine-tuning process on the first of the ten folds for configuration 1 and 41. In both instances, model learning for ATE plateaued before the third training step. In the third training step, AEC fine-tuning was performed in conjunction with ATE fine-tuning. With the hyperparameter settings in configuration 1, model learning for AEC remained minimal, starting at an F1-score of 38 and peaking at 42. With the settings in configuration 41, the model progressed from 38 to 49.
(a) F1-score per Epoch for Configuration 1.
(b) F1-score per Epoch for Configuration 41.
Figure 4: Progression of model performance during 10-fold cross validation.
Table 6: 10-Fold Cross Validation for Configuration 1 (Original Settings by [2]) and 41 (Highest Performance). The configuration with the highest performance is shown in bold.
Configuration 1
Cross Fold 1 2 3 4 5 6 7 8 9 10 Average
ATE 63.8 64.0 65.7 69.3 67.0 65.3 65.3 66.2 64.4 67.4 65.8
ATE+AEC 41.8 38.4 41.9 43.1 44.3 46.4 40.6 41.5 40.1 43.0 42.1
Configuration 41
ATE 70.3 67.9 71.2 71.5 69.0 64.5 71.4 69.8 63.6 70.9 69.1
ATE+AEC 49.3 45.9 52.3 49.6 46.7 47.7 47.2 47.7 44.1 49.0 48.0
5 Discussion
5.1 Training Dataset Annotation
The labeled ABEA dataset was created by human annotators through a process that aimed to balance strict, well-defined rules with the flexibility needed for the naturally ambiguous interpretation of emotions. Differences between theoretical conceptualization and practical applicability of the emotion classes and aspect terms warrant further discussion, as do the dataset results regarding their suitability for NLP model training.
5.1.1 Emotion ambiguity
The inherent subjectivity of human emotions, combined with the lack of contextual information such as facial expressions, spoken tone, or images, makes aspect-level labeling of emotions in written text a considerable challenge. Throughout the annotation process, many Tweets were encountered that could reasonably be labeled with two or more emotions, depending on the Tweet’s unknown context, to which none of the annotators had access. Despite the clear and logically sound structure of the emotion graph proposed by Shaver et al. [1], annotators repeatedly struggled to make definitive decisions on the emotion labels when encountering instances of sarcasm, rhetorical questions, or hypothetical statements. In many cases, sarcasm is self-evident to humans when provided with visual and auditory context. Even in writing, sarcasm can often be detected based through exaggerated emotional expression. Aside from such clear-cut cases, some Tweets fell into the fuzzy boundaries between earnest emotion and sarcasm, posing a challenge for the annotators to confidently label both the emotion and the aspect term. Similarly challenging were hypothetical scenarios, such as “if the Lakers don’t win this game, I might smash something fr”. These types of statements often involved conditional structures that juxtaposed different possible outcomes, each tied to a different emotion. Among other persistent challenges in distinguishing emotions, distinguishing between caring (Happiness) and worry (Fear), or frustration (Anger) and disappointment (Sadness) was frequently debated as conceptually overlapping emotions. The group annotation system and majority vote evaluation helped establish a consensus on these ambiguous cases.
Following the annotation process, the necessity of explicitly accounting for sarcasm and hypothetical statements stands out as a critical area for improvement. Future annotation guidelines may benefit from introducing distinct categories for these occurrences or excluding them from the dataset altogether. By addressing these edge cases directly, annotators could avoid ambiguity in their labels, ensuring greater consistency and improving the overall quality of the training data for emotion recognition models.
5.1.2 None Emotion Class
To ensure model consistency and reduce fine-tuning bias, a None class was added to account for neutral, unemotional statements. From an emotion labeling perspective, assigning None to an aspect term is counterintuitive, since an aspect term is by definition the explicit target of an emotion in a text. The class was nonetheless used to capture instances of neutral statements, without which an NLP model might force unemotional texts into one of the emotion classes, leading to bias and misrepresentations.
During initial experiments using the original GRACE hyperparameter settings, removing the None class led to a nearly 10 percentage-point decrease in ATE F1-scores. This drop indicates that, without a None class, the model struggled to systematically identify aspect terms. While the model became more robust during training with the None class, its inclusion introduced ambiguity into aspect term annotation. Since aspect terms tied to the None class did not logically serve as targets of emotions, the model may have struggled to transition between emotion and non-emotion contexts.
5.1.3 NLP-Appropriate Aspect Terms
Due to the nuanced and subjective nature of emotional expression, the conceptualization of what constitutes an aspect term further added to labeling uncertainties. The guidelines therefore established a strict definition of aspect terms, emphasizing nouns, pronouns, and verbs as suitable targets while largely excluding adjectives, articles, and descriptive phrases. This restriction was aimed at making the aspect terms consistent and easier for a model to learn through predictable patterns.
However, this computational focus diverges from the more nuanced human perception of what constitutes an aspect term. In reality, emotions are often directed at broader or more complex targets, extending beyond the central noun in a text. For example, in a Tweet such as "The decisions made by local officials have been disappointing," the emotional target includes the entire phrase "decisions made by local officials" rather than just "decisions" or "officials” on their own. Despite this discrepancy, limiting the scope of aspect terms was essential for systematic model learning, which relies on consistent input structures.
To test the effect of a less restrictive aspect term definition, a revised version of the dataset was created with a looser annotation of emotional targets. The preliminary results (cf. Figure 6) indicated that the model had significantly more difficulty extracting aspect terms, only reaching an F1-score of 46% with the original hyperparameter configuration, compared to the F1-score of 65% with the stricter definition. These results highlighted the need to define aspect terms in a way that balances model learnability while maintaining accurate representations of emotion targets.
5.1.4 Inter-annotator agreement
To achieve high IAA, the annotation process incorporated several strategies to enhance consistency among annotators. The annotators attended in-person sessions [10, 40], enabling discussions to resolve ambiguities as they progressed through the dataset. The effectiveness of this annotation approach was reflected by the fact that 89% of all aspect term annotations were included in the final dataset, based on the criteria of being labeled by at least two out of three annotators. This statistic was calculated at the character-level, i.e., the 11% of excluded characters also contained individually annotated whitespaces around aspect term sequences, or additional words adjacent to the final aspect terms. For the emotion labels, each identified aspect term sequence was similarly evaluated for a dominant emotion, which was identified in all but 12 cases that were individually reviewed. This high IAA, which enabled the majority of the initial dataset to be included in the final ABEA training dataset, underscores the value of annotator discussions.
From a social perspective, however, this methodological choice requires critical examination. Group dynamics, especially if pre-existing relationships or authority figures are involved, can bias annotators’ objectivity. Peer conformity theory [75, 76] suggests that individuals may align with dominant voices in group settings to avoid conflict, which can result in “groupthink”, biased labels, or loss of diverse interpretations. To counteract this issue, annotators were repeatedly assured that their own interpretations were valued and that no single person’s opinion was inherently more correct than any other. Any residual effects of group dynamics cannot be quantified precisely and must be accepted as a baseline risk in collaborative annotation.
5.2 Model training
5.2.1 Model performance plateau
Using the ABEA training dataset, the GRACE model was fine-tuned based on the highest average validation scores in the HPO configuration. Nevertheless, the training and validation results for ABEA consistently showed a clear upper boundary in terms of its performance on the joint ATE and AEC tasks. The 10-fold cross-validation yielded an F1-score of 48.0% for the joint task, while the glsATE task alone achieved 69.1%. By comparison, Luo et al. [2] reported 58.3% for the joint tasks and 75.7% for the glsATE-only task on the ABSA Twitter dataset from Mitchell et al. [35]. On the surface, these results, i.e., 10.3 and 5.8 point drops in comparative performance, can be intuitively attributed to the increased complexities of both the new dataset and the ABEA task. The training dataset consisted of more diverse word classes for aspect terms and was less domain-bound than conventional ABSA training datasets (cf. Figure 7). In addition, the identification of emotions rather than sentiment polarities introduced more granularity. Although these changes increase model learning difficulty and can reasonably be expected to cause a drop in performance relative to the ABSA baseline, a closer inspection of the training results provides deeper insights into the underlying causes for the model’s evident performance plateau.
Following the HPO, the cumulative results revealed that the GRACE model was not learning generalizable patterns for the two ABEA tasks, with training performance logs indicating a strong tendency towards overfitting on the training data. This tendency was apparent from a substantial difference between training and validation performance and could be traced back to the original hyperparameters and training settings reported by Luo et al. [2] as optimal for the ATE and ASC tasks, which also suffered from significant overfitting (cf. Figure 8). To reduce overfitting, several configurations were tested during HPO. Most effective among these were the increased dropout rates in BERT ’s hidden layers and the attention heads. At a dropout rate of 0.3, i.e., any given node connection in its network layers with 30% probability is dropped, the model’s training performance plateaued at F1-scores of 73% for the ATE task only and 61% for the joint ATE and AEC tasks, respectively. While this level of dropout successfully prevented overfitting, it did not lead to a more robust performance. Instead, validation metrics significantly dropped, indicating signs of underfitting, as reflected in low performance across both training and validation data.
Figure 5: Left: Validation metrics per epoch, Right: Training metrics per epoch during HPO when trying to adjust for overfitting (here showing hyperparameter configuration 3)
Consequently, the primary challenge during HPO was to strike a balance between avoiding underfitting and preventing overfitting. The optimization process highlighted this delicate trade-off, where reducing overfitting led to notable drops in both training and validation performance. This swing from overfitting directly into underfitting was evident in the validation metrics shown in Figure 5, where performance failed to improve and instead plateaued at relatively low F1-scores of 0.64 for ATE and 0.39 for AEC. The direct trade-off between decreased validation performance with regularization and decreased training performance without regularization implied that the GRACE model struggled to generalize patterns from the training data, regardless of regularization.
In light of this evident performance plateau, a critical examination of the training dataset itself is essential Since measures to ensure label consistency were taken, the weakness is therefore likely connected to data quantity [77]. Mosbach et al. [78] argue that BERT fine-tuning instability is not caused by small training datasets themselves, but rather by the reduction of training iterations which often accompany the use of smaller datasets. However, in this study, several settings for increased training iterations did not yield significant performance improvements. A small dataset like the 2,621-instance ABEA dataset likely hinders model learning for complex tasks by failing to provide a sufficiently diverse range of examples, resulting in narrow representations that do not capture the variability needed for real-world application. This can lead to overfitting, where the model memorizes specific patterns in the training dataset, or underfitting when regularization is applied due to a lack of representative data.
The impact of a too small dataset is further supported when considering ATE and AEC performances. Across all HPO configurations, the performance for ATE was considerably higher than for ATE and AEC combined. A likely reason for the higher ATE performance is the amount of training examples available for aspect terms, as it includes all 5,419 terms across emotion classes. By comparison, for each emotion class, far fewer examples were available, e.g., only 282 for the class Fear. The ATE task therefore benefited from a more diverse training set, even with the added complexity of more varied word types compared to prior ABSA datasets. In contrast, the number of training examples per emotion class was much lower, especially for the classes Fear and Sadness. Additionally, many words were unique to a single emotion class. While this may be expected for a small dataset, it increases the risk of the model learning specific words rather than generalizable patterns. In future work, data augmentation techniques should be considered to expand small training datasets without requiring further time-consuming annotation. This would be a viable option to mitigate overfitting and help NLP models to generalize learned patterns.
5.2.2 Hyperparameter settings
Throughout the HPO process, 16 hyperparameter categories were adjusted. Some settings, such as weight decay, layer reduction, and layer freezing, were aimed at mitigating overfitting. Other modifications were intended to improve learning capacity, such as incorporating more complex classifier architectures or introducing earlier split between the ATE and AEC branches. Overall, the average model performance was relatively insensitive to changes in hyperparameters, with only minor fluctuations across the different HPO configurations. This suggests that insufficient dataset size, rather than hyperparameter settings, may be the primary limiting factor in model performance. Nonetheless, some hyperparameters adjustments did yield noticeable effects on model outcomes.
One of the most influential changes was the adjustment of the classification heads. Introducing more complex classification head architectures, specifically DenseNet-inspired classifiers, led to slight performance gains in the AEC task. However, this improvement came with a significant trade-off for the ATE task, particularly on the SemEval validation datasets, ultimately reducing overall performance substantially. The increased complexity of the classification heads likely exceeded what the dataset size could support, causing more severe overfitting on the training data. To balance training and validation performance, the final model retained a linear layer as classification head.
VAT was originally included in the training process of GRACE during the second training step. However, HPO results indicated that using VAT had no positive impact on model performance. One plausible explanation is that the ABEA training data is complex enough without additional adversarial examples, making VAT unnecessary or even detrimental. As a result, VAT was removed from the final model configuration to streamline the training process. For an augmented dataset, however, VAT should be re-considered and may achieve its intended effect of increasing model robustness.
Various GHL configurations were tested, including its application at different training steps, adjustments to the number of bins (used to group and assign weights to training examples based on their difficulty), and modifications to momentum values. While increasing the number of bins from 24 to 32 yielded the highest average performance by allowing finer differentiation of training examples based on difficulty, none of the GHL adaptations resulted in a significant improvement over previously achieved performances. As a result, the original GHL settings were retained. However, the underlying assumption of the GHL mechanism should be critically evaluated in and of itself. GHL assumes that the most frequent labels are also the easiest to learn. However, this may not apply to emotion classification, where the class distribution does not necessarily correlate with learning difficulty. In the new ABEA dataset, it is not at all self-evident that the most frequent classes, Happiness and None, are easier to learn compared to the least frequent classes Fear and Sadness. In contrast, for the original ABSA dataset [35], GHL is far more applicable since 70% of all examples belong to the Neutral sentiment class.
Adjustments to the number of shared layers between the ATE and AEC branches, as well as the number of decoder layers in the AEC branch, led to minor improvements in average performance. Originally, the GRACE model used nine shared layers, with only two decoder layers in the AEC branch. Interestingly, experiments showed that reducing the number of shared layers to five and accordingly increasing the AEC branch’s decoder layers to six improved model performance. This suggests that the model was already learning the ATE task satisfactorily, allowing the additional dedicated layers in the AEC branch to enhance its ability to capture task-specific patterns.
Layer freezing was tested as a regularization strategy to mitigate overfitting. Configurations included freezing the bottom three, five, and nine layers of BERT-base’s twelve overall layers. Surprisingly, freezing the bottom three and five layers had minimal impact on the model’s performance, either positively or negatively, with the averaged performance remaining at 49.4% and 49.5%, respectively. Freezing the bottom nine layers, however, did lead to a noticeable performance drop to an average of 45.1%. This finding further reinforced the notion that hyperparameter tuning had limited effect and suggested the dataset constraints, rather than the model structure, as the primary bottleneck for performance gains.
6 Conclusion
This paper addresses existing research gaps in the field of Aspect-based Emotion Analysis by creating a consistent training dataset (RQ1) as well as adapting and fine-tuning the GRACE model [2] for the joint extraction of the sub-tasks ATE and AEC (RQ2).
To address the training dataset bottleneck for ABEA, a Twitter dataset was labeled using group annotation and majority voting to ensure label consistency. In particular, the introduction of a None emotion class and a tightening of aspect term definitions addressed the conceptual ambiguities inherent to emotion-classification and NLP at the token-level, leading to improved annotation coherence. The final dataset comprised 2,621 Tweets.
To repurpose the ABSA model GRACE [2] for ABEA, extensive HPO was conducted. The most impactful hyperparameter changes on model performance included reductions in batch size, a warm-up method adjustment, a reduction in the number of shared layers between the ATE and AEC branches, an increase of AEC branch layers, extending training epochs, and omitting VAT in the second training step. A more complex classification layer negatively impacted model performance. The fine-tuned ABEA model plateaued at F1-scores of 70.1% for ATE and 46.9% for joint ATE and AEC.
The lower performance of EmoGRACE for ABEA compared to the reported performance of GRACE for ABSA can be largely attributed to three factors. First, the shift from sentiment polarity to more complex emotion classification. Second, the increased variance in types of words for aspect terms, extending beyond just nouns. Third, the size of the training dataset proved insufficient to allow for generalizable model learning at the high level of task complexity, indicating that data augmentation may be a necessary pre-requisite for continued ABEA research. Future work may consider the use of generative foundation models, such as Generative Pre-Trained Transformer (GPT) models [79], which have shown strong capabilities in handling a variety of NLP tasks and may even outperform fine-tuned task-specific BERT models in certain applications [80].
The effectiveness of the presented ABEA method in capturing public emotions should be evaluated through a targeted, domain-specific case study, such as analyzing posts during a specific natural disaster. The model could be applied to the social media-based analysis of other highly emotional social topics, such as public health crises [81], climate change [82], urban events [59] or political discourses [83], providing insights into collective emotions and their targets.
[…]
Acknowledgements
This project has received funding from the European Commission - European Union under HORIZON EUROPE (HORIZON Research and Innovation Actions) under grant agreement 101093003 (HORIZON-CL4-2022-DATA-01-01). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union - European Commission. Neither the European Commission nor the European Union can be held responsible for them.
This paper also received funding through the Austrian Research Promotion Agency (FFG) and the DigitalInnovationLayer project (FO999898902).
We would like to thank David Hanny and Shaily Gandhi (IT:U) for their feedback.
[…]
Author contributions Conceptualization, C.Z., B.R.; methodology, C.Z.; software, C.Z.; validation, C.Z.; formal analysis, C.Z.; investigation, C.Z.; resources, B.R.; data curation, C.Z., S.S.; writing—original draft preparation, C.Z., S.S.; writing—review and editing, C.Z., S.S.; visualization, C.Z.; supervision, B.R.; project administration, B.R.; funding acquisition, B.R.
All authors have read and agreed to the published version of the manuscript.
Appendix A
Table 7: Training datasets found in relation to the search terms “aspect-based emotion analysis”, “aspect-based emotion dataset”, or “fine-grained emotion dataset”.
Dataset Public Data Source Size Comment
[11] Yes Adidas Instagram account (images and comments) 4,900 Multimodal dataset
[43] No Twitter 5,000 Dataset and model serve the understanding of financial/stock market risk and opportunity
[40] No SemEval Restaurant dataset [31] 3,850 Multilingual, mostly Indonesian language; After IAA the authors removed emotions ‘anger’, ‘fear’ and ‘disgust’ due to small cohorts
[26] No Indonesian restaurant review platform ‘PergiKuliner’ 14,103 Multilingual (English, Dutch, German, French); designed for a business platform
[41] No Bol.com, Trustpilot, TripAdvisor, and proprietary data from project partners 1,000+ Hungarian language
[42] No Hungarian speeches from 2014 - 2018 1,008 Aspects are not labeled per text, but considered as overall categories for which emotion scores are identified; emotions are based on COVID-19 death rates
[84] No Twitter unclear Aspects are not labeled per text, but considered as overall categories for which emotion scores are identified; emotions are based on COVID-19 death rates
Figure 6: Training metrics on an adapted training dataset with more loosely defined aspect terms.
Figure 7: Direct comparison of the type of words labeled as aspect terms among ABSA and ABEA datasets.
Figure 8: Left: Validation metrics per epoch, Right: Training metrics per epoch using the original ABSA dataset with the original hyperparameters and training settings as reported by Luo et al. [2]